# Towards Generalist Robot Policies: What Matters in Building Vision-Language-Action Models

### v3

@@ I Introduction
| Q2: Which Backbone | Q2.1: Which type of VLMs are more suitable for constructing VLAs? | A2.1: Sufficient vision-language pre-trained on large vision-language datasets benefit VLAs |
| Q3: How to Formulate | Q3.1 : What is the best-performing VLA structure? | A3.1: Continuous action space with policy head to integrate history is the best structure. |
| Q3.2: How do different formulations affect the generalization and data efficiency for VLAs? | A3.2: The KosMos backbone with a separate policy head for history fusion performs the best in generalization and data efficiency. | |
| Q4: When to Leverage Extra Data | Q4.1: How do large-scale cross-embodiment datasets contribute to VLAs? | A4.1: Extra in-domain data shows beneficial; 2) Post-training further improves overall as well as few-shot performance. |
II Main Results and Findings
Vision-language-action models (VLAs) are commonly defined as models fine-tuned from pre-trained large-scale vision-language models (VLMs) using imitation learning [24, 7]. By leveraging the robust vision-language representation capabilities of VLMs, VLAs offer a promising approach for developing generalist robotic policies capable of handling complex tasks. However, this approach is not universally accepted as the sole or optimal solution. For instance, modular approaches utilize pre-trained vision and language modules to encode latent representations of multi-modal inputs [31, 6], while alternative methods rely on direct training with diverse robotic datasets [39]. Even within VLA research, there is no consensus on architectures or training recipes [24, 8, 22, 7].
The primary objective of this work is to establish VLAs as robust generalist robotic policies by thoroughly analyzing contemporary VLA architectures and identifying the key factors driving their performance. To this end, we introduce RoboVLMs, a unified VLA framework that enables the systematic integration and exploration of essential components in VLA design. Using this framework, we conduct extensive experiments to address several critical questions.
1.
Why do we prefer VLAs?
2.
How should we formulate VLAs?
3.
Which VLM backbone is more suitable for VLAs?
4.
When should we leverage cross-embodiment datasets?
As shown in Tab. I, we further divide the 4 essential problems into 6 research problem, and implement successive experiments of VLAs to answer each research problem. The built VLA model with a proper backbone and structure can outperform the state-of-the-art generalist robot policies by a large margin. We hope that our findings can practically help build robust, generalizable, and well-performing VLAs.
To comprehensively evaluate the performance of VLAs, in this work, we benchmark all models on a diverse set of benchmarks and robotic manipulation tasks in both simulation and the real world. Specifically, as shown in Fig. 3, we choose tow well-known and widely used simulation benchmarks (CALVIN [32] and SimplerEnv [40]), and a real-world robot manipulation experiment to evaluate VLA models:
CALVIN [32] is a simulation benchmark for multitask table-top manipulation. The dataset contains four splits A, B, C, and D according to different scene settings and provides 34 basic tasks with 24K human teleoperated demonstrations annotated with language instructions in total. Evaluation metrics include the success rates of finishing consecutive tasks, as well as the average number of achieved tasks successfully executed (shorted as Avg. Len.).
SimplerEnv [25] is designed as a suite of real-to-sim environments and enables the evaluation of robot policies in simulation. It creates a comparable arena for benchmarking the success rate of robot policies in private real-world settings such as Google Robot [6, 7] and Bridge V2 [45].
Real Robot Benchmark [8] consists of over 70K teleoperated human trajectories used to fine-tune robot policies, covering 105 manipulation tasks. To evaluate the performance of models on this benchmark, we adopt the approach outlined in [23], testing each model on one Simple setting and four challenging Unseen settings. Examples of these settings are shown in Fig.4. In total, we evaluate each VLA across 20 tasks, with 5 settings per task and 3 rollouts per setting, reporting the average success rate for each setting. A detailed description of the benchmarks is provided in Appendix K and Appendix D.
All tasks included in these benchmarks are driven by single-arm robots, leading to a 7-DoF action - the 6D pose 222Represented in Euler angles. of the gripper and one-dimensional open/close status. Robot observation is accessible from proprioceptive sensory information, visual observation, and language input.
II-A Why do we prefer VLAs?
@@
This section investigates one of the key questions: Why do we want VLAs? To answer this question, we need to answer the following prerequisite sub-question at first:
Concretely, we demonstrate the best-performing VLA resulting from our study, which sets a new state-of-the-art result on both CALVIN and SimplerEnv benchmarks, outperforming all other robot policies with a clear margin. All results are shown in Tab. II and Fig. 5.
From these tables, we can see that our strongest RoboVLM exceeds the existing state-of-the-art generalist policies by a large margin and establishes a strong baseline for robot manipulation tasks both in simulation and real-world experiments. Concretely, we can easily observe the following facts:
On CALVIN benchmark, our best model achieves the highest performance in all metrics and demonstrates superior generalization ability when transferring from ABC to D (a novel scene unseen in the training splits) with an absolute improvement of 12.6% for the execution of a single task and a total improvement of 30.3% for 5 consecutive tasks. On average, under zero-shot settings, our model can finish 4.25 tasks out of 5 tasks for each single rollout, outperforming the previous SOTA model (GR-1) by 1.09 tasks.
We also investigated the impact of vision-language pre-training on the generalization and data efficiently (Fig. 6 and Tab. IV), and the detailed result in shown in Appendix H. For the generalization in CALVIN, we adopt the official setting: training models on the split of ABC and validating performance on D. To evaluate data efficiency, we conduct experiments on model scales ranging from 3B to 9B and various data scales: 10% training data (0.1x ABCD) the standard setting (ABCD), and 500% training data (5x ABCD). The additional data originates from the officially released unlabeled datasets, following the setups as introduced in Wu et al. [47], the detailed results on different data scales are shown in Appendix E.
As discussed above, we deploy the best-performing RoboVLM model, that is, the one based on the decoder-only KosMos in real-world scenarios to validate its effectiveness. As shown in Fig. 4, our experiment involves 20 tasks with multiple skills, including Open, Close, Press Button, Pick & Place, etc. For each task, we evaluate five rollouts, with the basic setting, novel skill description, unseen distractors, unseen target object, and unseen background.
Our robot system for real experiments is built on a 7-DoF Kinova Gen3 robot arm paired with a Robotiq 2F-85 gripper, please refer to Sec. IV for more details of the real robot. For input, we take the RGB images for the two cameras equipped on the robot head and wrist separately. The head camera provides an overview of the workspace while the gripper camera offers a close observation of the interaction area between the end effector and the environment.
We fine-tune Octo-Base, OpenVLA, and KosMos P.H. built by RoboVLMs on the real robot benchmark and compare their performance. The result is shown in Fig. 7. We observe that the best VLA (KosMos P.H.) built by RoboVLMs achieves the best performance in all evaluation setups, extremely on Simple and Unseen Background, demonstrating their effectiveness and generalization ability, which is consistent with the results in SimplerEnv and CALVIN simulation.
The qualitative results are shown in Appendix K, including success rollouts in various settings and some representative failure cases. KosMos P.H. not only outperforms baseline models in basic setting tasks like Open Drawer, Pickup Egglant and Press the Toaster Switch, but also achieves better performance over unseen objects, distractors, and backgrounds. Furthermore, as shown in Fig. 8, KosMos P.H. emerges with self-correction ability, it can realize the incorrect positions of the end effector and correct its future trajectory to complete the task successfully. Note that this ability does not appear in the other tested baselines, and this kind of data is not contained in the training dataset.
II-B How should we formulate VLAs?
In this section, our study addresses questions regarding VLA formulations, including different design choices of VLA structures and various backbones. To answer these questions, we conduct a series of controlled experimental studies to ablate various VLA formulations on the CALVIN benchmark for rapid evaluation.
| 1 | 2 | 3 | 4 | 5 | ||||
| LLaVA | One-Step | Disc. | 0.809 | 0.484 | 0.278 | 0.175 | 0.103 | 1.85 |
| One-Step | Cont. | 0.793 | 0.592 | 0.420 | 0.329 | 0.235 | 2.37 | |
| Interleaved | Cont. | 0.892 | 0.645 | 0.436 | 0.282 | 0.181 | 2.44 | |
| Policy-Head | Cont. | 0.873 | 0.678 | 0.506 | 0.376 | 0.275 | 2.71 | |
| Flamingo | One-Step | Disc. | 0.681 | 0.318 | 0.133 | 0.062 | 0.029 | 1.22 |
| One-Step | Cont. | 0.681 | 0.354 | 0.158 | 0.076 | 0.035 | 1.30 | |
| Policy-Head | Cont. | 0.964 | 0.896 | 0.824 | 0.740 | 0.662 | 4.09 | |
| KosMos | One-Step | Disc. | 0.424 | 0.097 | 0.023 | 0.005 | 0.002 | 0.55 |
| One-Step | Cont. | 0.881 | 0.599 | 0.364 | 0.221 | 0.124 | 2.19 | |
| Interleaved | Cont. | 0.987 | 0.915 | 0.824 | 0.737 | 0.660 | 4.12 | |
| Policy-Head | Cont. | 0.967 | 0.930 | 0.899 | 0.865 | 0.826 | 4.49 | |
The performance of various VLA structures in CALVIN is summarized in Tab. III. From these results, we draw the following key observations:
Continuous action matters: By comparing two types of action spaces, continuous and discrete, as shown in Tab. III, we observe that under the single-frame formulation, continuous action spaces consistently outperform discrete ones, particularly as task horizons increase. This finding is intuitive: continuous actions can represent high-precision floating-point values, whereas discrete actions are limited to indexing action intervals. For long-horizon tasks, the accumulation of compounding errors significantly degrades the performance of discrete actions.
History observation matters: As shown in Tab. III, under the same VLM structure (either Encoder-Decoder or Decoder-only), models incorporating history observations as input consistently outperform one-step models, achieving substantially higher success rates across all tasks. This improvement holds regardless of the history fusion strategy used. Furthermore, increasing the length of an observable history can enhance performance, albeit at the cost of higher computational overhead.
@@ II-B How should we formulate VLAs?
However, beyond the performance itself, one of the most important challenges for modern VLAs is achieving generalization to novel objects and environmental settings, which is critical for practical deployment across various robots and scenarios. Conversely, when generalization is insufficient, fine-tuning the policy with a few new demonstrations becomes ideal. Thus, VLAs should inherit the generalization capabilities of VLMs in open-world settings while maintaining high data efficiency when additional in-domain training samples are available. Therefore, we further investigate the following question:
To address the question, we empirically study and evaluate the generalization and data efficiency of various VLA formulations, aiming to provide practical insights for training high-performing VLAs. Specifically, we assess the generalization and data efficiency of different VLAs built with RoboVLMs by training models with different architectures and formulations on varying data scales using the CALVIN datasets. As discussed earlier, we focus on comparing the interleaved and policy head formulations using the OpenFlamingo and KosMos backbones, which have shown strong potential among all configurations. Note that the interleaved formulation can only be paired with a decoder-only structure. The results presented in Fig. 9 and Tab. IV, lead to the following observations:
For generalization performance (Fig. 9), our best model, based on the KosMos backbone and leveraging a policy head for history fusion, exhibits only a slight performance drop in zero-shot settings. In contrast, other formulations experience significant performance declines. This finding highlights that the model architecture significantly impacts generalization. This conclusion is further supported by results in Fig. 5, where tasks in the evaluation set are paired with novel instructions, and Fig. 7, where our best model outperforms others by a large margin across all unseen tasks.
II-C Which VLM backbone is better for VLAs?
@@
Following such a finding, the choice of using a policy head for history fusion in VLA can be finalized. However, a critical question remains when selecting the most appropriate VLM to build our VLA:
To thoroughly investigate this question, it would be ideal to conduct experiments in highly controlled settings. However, training VLMs on large-scale vision-language datasets is extremely resource-intensive. Therefore, we base our VLAs on a diverse selection of pre-trained large-scale vision-language backbones with varying architectures, training data scales, model sizes, and latent embeddings. These include Flamingo model family [1] (Encoder-Decoder), and a series of decoder-only VLMs, including LLaVA [28], Qwen-VL [2], MoonDream [44], UForm [41], Paligemma [3], and KosMos [36]. Noticeably, in this section, for fair comparisons, all the models are trained with static images instead of both static and hand cameras.
Although this approach may not offer a fully controlled comparison, our extensive experiments aim to provide insights into the impact of different VLM backbones on VLAs. The results, presented in Tab. V, reveal the following observations:
KosMos and Paligemma demonstrate the distinctively better performance: From Tab. V, we can see that these two backbones are much better than others with a significantly clear margin. Their superior performance benefits from sufficient vision-language pre-trained on large vision-language datasets. This outcome is intuitive, as extensive pre-training facilitates stronger alignment between visual and linguistic features—an alignment critical for language-conditioned manipulation tasks.
We discuss more influencing factors and interesting findings in Sec. III.
II-D When should we leverage cross-embodiment datasets?
To address this, we break the question into two sub-questions:
1.
2.
When and how should these data be utilized effectively?
In this section, we conduct a series of experiments to investigate different strategies for using external large-scale cross-embodiment datasets. Specifically, we explore two primary settings:
Post-train: First, training the VLMs on cross-embodiment datasets, followed by fine-tuning with in-domain manipulation tasks. This strategy has been adopted by [4].
For comparison, we also evaluate a baseline setting, Finetune, where the VLA is trained exclusively on in-domain data. Additionally, for Google Robot, we include both RT Partial Finetune and RT Finetune, where RT Partial Finetune involves only the trajectories with the same task type as the evaluating tasks, and RT Finetune involves co-finetuning the policy with additional data from the same robot across different tasks. For Bridge, we only evaluate Bridge Finetune which finetune the policy with the entire Bridge-V2 dataset, since the training dataset does not contain trajectories with the same instructions of the evaluated tasks.
To evaluate the impact of cross-embodiment datasets more comprehensively, we also perform experiments on CALVIN, which is not part of OXE. For CALVIN, we omit the co-train setting and mainly focus on whether cross-embodiment datasets benefit few-shot learning for robot manipulation on out-of-distribution tasks. Therefore, we conduct experiments on CALVIN using only 10 trajectories per task (CALVIN few-shot). To keep settings consistent, we only take the image from the static on-head camera as input.
The comparison of leveraging cross embodiment at different training stage is shown in Fig. 10. From up to down, the figure shows the evaluation results for SimplerEnv-Google Robot and SimplerEnv-Bridge. Further, the comparison for whether to integrate cross embodiment pre-train on CALVIN few-shot is shown in Fig. 11. The following observations can be drawn:
Pre-training with cross-embodiment data does not help significantly. Comparing OXE Pre-train and RT-Partial Finetune reveals that for both Google Robot and Bridge, co-training with cross-embodiment data does not lead to substantial performance improvements. In particular, for Google Robot, training with additional in-domain data (RT Finetune)—even from different tasks—achieves higher success rates (compared with RT-Partial Finetune). This indicates that in-domain data, even if task-agnostic, are more effective for improving model performance than cross-embodiment data.
Post-training after cross-embodiment pre-training shows potential benefits. The average performance of the post-trained model (52% on Google Robot and 38% on Bridge) exceeds that of the model fine-tuned exclusively on in-domain data (48% on Google Robot and 31% on Bridge). This suggests that cross-embodiment pre-training can provide a useful initialization that benefits subsequent fine-tuning.
Pre-training improves few-shot learning performance. In the few-shot setting for CALVIN, with a single-view on-head camera, pre-training significantly improves the performance by 17.2% in terms of single-task execution and 0.25 more executed tasks in each rollout. It can be concluded that pre-training on large-scale cross-embodiment datasets benefits the learning of more effective representations for robot manipulation, which can quickly adapt to novel manipulation tasks with unseen objects and environmental settings.
III Discussion
@@
Under investigated observations. During our experiments, we found that VLAs built upon Qwen-VL and LLaVA, the performance is surprisingly low, compared with their original performance on vision-language tasks. After adding a perceiver resampler after the vision encoder, we found that the VLAs based on Qwen-VL and LLaVA could obtain great performance gain and reach reasonable performance. We hypothesize that the performance gain is related to the image resolution and number of vision tokens in the input token sequence.
Limitations. Although we make every effort to investigate the key challenges in building Vision-Language Agents (VLAs), this work remains preliminary and has several limitations at the current stage. (1) For the sake of quick and simple expansion over existing Vision-Language Models (VLMs), we retain the multi-modal interaction structure within the VLM (e.g., attention masks, mixture of experts). On top of this, we further develop the interaction between vision, language, and actions, which is a common approach in most existing works [22, 24]. However, a specialized design for the architecture and multi-modal interaction with actions has the potential to yield superior performance (e.g., the model [4]), and warrants further exploration. (2) The categorizations and formulations of VLAs considered here are simplified and limited for the reasons outlined in (1). (3) The action tokenization, policy head, and corresponding training objectives are not fully explored in this work. For example, techniques like VQ-VAE [42], diffusion models [17, 9], and flow matching [26, 12, 4] remain under-explored in the context of VLAs. (4) The set of VLM backbones considered in this study is limited and can be actively expanded. (5) Deploying such large models for real-time robotic control remains a significant challenge.
Future works. For future work, we envision several potential directions for advancing generalist robot policies. 1) As aforementioned, our current approach faces limitations in the design of internal structures for VLMs, policy heads, and corresponding training objectives. Further investigation into more fine-grained design choices for VLAs could be highly valuable, as recent studies suggest they play a significant role in improving efficiency and effectiveness [4]. 2) Beyond semantic generalization, an ideal generalist robot policy should be capable of handling long-horizon, complex task instructions (e.g., make breakfast), reasoning through executable actions step by step, and generating meaningful physical interactions with its environment (e.g., [52]). In our future work, we aim to explore the key elements required to develop policies with these advanced capabilities.
@@
IV Method and Material
We consider the problem of controlling a robot to accomplish a set of tasks, based on language instructions , and historical observations at every time step with the maximum historical length . In this paper, we mainly consider a table-top robot arm, therefore the observations are sensor inputs and images, e.g., , from either a third-view camera, a gripper camera, or both. We build a controlling policy , where the action is modeled as a 7-dim vector, including the 6-DoF pose of the gripper, along with its open/close status. In this section, we will first introduce the formulation of vision-language models (Section IV-A), based on which we build our VLA formulations.
IV-A Vision Language Model
Decoder-only. Decoder-only architectures, in contrast, rely on a unified transformer framework where both the input modalities (vision and text) and the output sequences are processed in the same auto-regressive decoder. In these models, visual features are first embedded into token-like representations (via vision processors), which are then concatenated with textual tokens and passed through the decoder. Multi-modal feature fusion occurs naturally through the self-attention mechanism, allowing the decoder to model dependencies between visual and textual inputs during token generation. Decoder-only architectures are more flexible and scalable, making them well-suited for tasks like instruction following, multi-modal question answering, and open-ended generation. Examples of decoder-only models include GPT-4V [49] and LLaVA [28].
IV-B Vision-Language-Action Models
where are a sequence of predicted 7-dim actions, is the action sequence length and is the history observation length. Different from VLMs, the observations of VLAs usually contain proprioceptive states like the joint angles and end-effector positions besides the visual inputs . As is discussed in Section IV-B, we abstract and categorized VLAs into four representative structures based on 1) historical information modeling and 2) action space. Before we describe these distinct models in form, we first introduce the general pre-process and prediction principles when dealing with robot actions in different action spaces.
IV-B1 Action Pre-process
is the normalized action, each dimension is in the range of , the last dimension representing the open/close status of the gripper . During inference time, we will reversely map the predicated actions into unnormalized actions.
IV-B2 Action Prediction
Continuous Actions. We optimize the mean square error (MSE) and binary cross entropy (BCE) for the predicted action sequence with the ground truth action sequence :
where the MSE loss is computed for the first six dimensions, and BCE loss is for the last gripper dimension between the predicted action and the ground truth , is the balancing weight.
IV-C VLA Structures
As shown in Fig. 12, there are mainly four kinds of VLA structures categorized by action space and history aggregating method, namely one-step-continuous-action models, one-step-discrete-action models, interleaved-continuous-action models, and policy-head-continuous-action models. Note that our proposed framework RoboVLMs could transfer VLMs to arbitrary VLA structures with no effort. We will illustrations the considered four VLA structures in detail in the following section.
IV-C1 One-step Models
One-step models predict future action sequences using only the observation at the current time step , i.e., a history length of 1.
| [LRN] | (10) | |||
Discrete-action model: For discrete action prediction, we directly follow the straightforward next-word prediction same as VLMs, where actions are discretized into tokens like texts:
IV-C2 Interleaved-Continuous-Action Models
Interleaved models receive observation-action sequences:
The which is utilized to predict the action chunk , represents the [LRN] inserted after and fused with the observations before . The loss and action unnormalization procedure is identity with one-step continuous action models. At time instant of inference, the input sequence contains only the current observation and the language instruction , we add the learnable token [ACT] at the end of the input sequence and pass the sequence to the VLM to predict the action. After the robot executes the predicted action, we add the new observation and language instruction to the input sequence to predict the action in the current step. The interleaved-continuous-action models include GR-1 [47], OCTO [39], GR-2 [8]. Note that the interleaved-discrete-action models like GATO [38] and RoboCat [5] are out of consideration.
IV-C3 Policy-Head-Continuous-Action Models
IV-D Real Robot Platform
As shown in Fig. 13, our real robot platform consists of a Kinova Gen-3 robot arm, equipped with a Robotiq 2F-85 parallel-jaw gripper and two cameras: one static camera for capturing the workspace and another camera mounted on the end-effector. The static camera is a Kinect Azure, while the wrist-mounted camera is a RealSense D435i. The workspace is a 55 cm x 24 cm table, and there are more than 40 objects distributed across the evaluated scenes.
IV-E Discussions about Structures
Appendix A Acknowledgments
We thank all the members of the robotics research team at ByteDance Research for their assistance in real-world data collection, setup design, robot maintenance, and experiments. The author Minghuan Liu is supported by the ByteDance Scholarship. We also want to thank @YouJiacheng for his active and instructive discussion on X.
Appendix B Contributions
Project Leads: Xinghang Li, Hanbo Zhang, Minghuan Liu
Methodology and Codebase: Xinghang Li
Model Training and Evaluation (experimental design, implementation): Xinghang Li, Hanbo Zhang, Dong Wang, Minghuan Liu, Xiao Ma, Jirong Liu
Real-Robot Deployment & Experiments: Xinghang Li, Peiyan Li
Paper (logic, figures, visualizations, writing): Xinghang Li, Minghuan Liu, Hanbo Zhang, Bingyi Kang, Xiao Ma, Peiyan Li, Jirong Liu, Huaping Liu, Tao Kong
Advising: Huaping Liu, Tao Kong, Hanbo Zhang, Xiao Ma, Bingyi Kang
Appendix C Implementation Details

### v4

@@ VII Contributions
| Q2: Which Backbone | Q2: Which type of VLMs are more suitable for constructing VLAs? | A2: Sufficient vision-language pre-trained on large vision-language datasets benefit VLAs |
| Q3: How to Formulate | Q3.1 : How to select the best-performing VLA structure? | A3.1: Continuous action space with policy head to integrate history is the best structure. |
| Q3.2: How do different formulations affect the generalization and data efficiency for VLAs? | A3.2: Leveraging policy head for history fusion is the best in terms of generalization and data efficiency. | |
| Q3.3: How to design appropriate training objective and execution paradigm? | A3.3: Diffusion and MSE training loss are comparable, executing action chunk is efficient and effective. | |
| Q3.4: Does Mix-of-Expert structure benefit VLA learning? | A3.4: Introducing the Mix-of-Expert structure can improve the performance and generalization of VLAs. | |
| Q4: When to Leverage Extra Data | Q4: Which training stage should introduce cross-embodiment datasets? | A4: 1) Extra in-domain data shows beneficial; 2) Post-training further improves high-frequency tasks as well as few-shot performance. |
@@
| RT-1-X | 0.167 | 0.000 | 0.208 | 0.042 | 0.083 | 0.000 | 0.000 | 0.000 |
| Octo-Base | 0.347 | 0.125 | 0.528 | 0.083 | 0.319 | 0.000 | 0.667 | 0.431 |
| Octo-Small | 0.778 | 0.472 | 0.278 | 0.097 | 0.403 | 0.042 | 0.875 | 0.569 |
| OpenVLA-7b | 0.041 | 0.000 | 0.333 | 0.000 | 0.125 | 0.000 | 0.083 | 0.041 |
| RoboVLMs (Ours) | 0.583 | 0.458 | 0.458 | 0.417 | 0.375 | 0.167 | 0.958 | 0.958 |
Standing | Average | Average | Open | Close | Average | Average | ||||
| RoboVLMs (Ours) | 1.000 | 0.910 | 1.000 | 0.970 | 0.888 | 0.565 | 0.537 | 0.551 | 0.325 | |||
| Cross-Emb Pre-Train | 0.458 | 0.458 | 0.333 | 0.250 | 0.458 | 0.042 | 0.916 | 0.833 |
| In-domain Full Finetune | 0.625 | 0.417 | 0.500 | 0.250 | 0.417 | 0.083 | 1.000 | 1.000 |
| Post Train | 0.583 | 0.458 | 0.458 | 0.417 | 0.375 | 0.167 | 0.958 | 0.958 |
Standing | Average | Average | Open | Close | Average | Average | ||||
| Cross-Emb Pre-Train | 0.940 | 0.840 | 0.890 | 0.890 | 0.571 | 0.222 | 0.444 | 0.333 | 0.019 | |||
| Target Task Finetune | 0.420 | 0.320 | 0.440 | 0.400 | 0.448 | 0.167 | 0.567 | 0.368 | 0.000 | |||
| In-domain Full Finetune | 1.000 | 0.910 | 1.000 | 0.970 | 0.888 | 0.565 | 0.537 | 0.551 | 0.325 | |||
| Post Train | 0.990 | 0.960 | 1.000 | 0.980 | 0.740 | 0.250 | 0.670 | 0.460 | 0.050 | |||
@@
Appendix A Discussion
@@
Under investigated observations. During our experiments, we found that VLAs built upon Qwen-VL and LLaVA, the performance is surprisingly low, compared with their original performance on vision-language tasks. After adding a perceiver resampler after the vision encoder, we found that the VLAs based on Qwen-VL and LLaVA could obtain great performance gains and reach reasonable performance. We hypothesize that the performance gain is related to the image resolution and the number of vision tokens in the input token sequence.
Limitations. Although we make every effort to investigate the key challenges in building Vision-Language Agents (VLAs), this work remains preliminary and has several limitations at the current stage. (1) The action tokenization, policy head, and corresponding training objectives are not fully explored in this work. For example, techniques like VQ-VAE, VQGAN, and FAST action tokenizer remain under-explored in the context of VLAs. (2) The set of VLM backbones considered in this study is limited and can be actively expanded. (3) Deploying such large models for real-time robotic control remains a significant challenge.
Future works. For future work, we envision several potential directions for advancing generalist robot policies. Beyond semantic generalization, an ideal generalist robot policy should be capable of handling long-horizon, complex task instructions (e.g., make breakfast), reasoning through executable actions step by step, and generating meaningful physical interactions with its environment. In our future work, we aim to explore the key elements required to develop policies with these advanced capabilities.
@@
Appendix B Implementation Details